OpenStudy (anonymous):

(log_(3) (log_(3) x)) = -1 solve for x

OpenStudy (jdoe0001):

\(\Large log_3[log_3(x)]=-1\quad ?\)

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

do you know how to convert log notation to exponential notation?

OpenStudy (anonymous):

kind of this one would be 3^-1 = 1/3 so log_3(1/3) ?

OpenStudy (anonymous):

because a^b=c and B^c=a

OpenStudy (jdoe0001):

well, more or less, yes \(\bf log_3[log_3(x)]=-1\implies 3^{-1}=log_3(x)\implies \cfrac{1}{3}=log_3(x)\\ \quad \\ \cfrac{1}{3}=log_3(x)\implies \large 3^{\frac{1}{3}}=x\)

OpenStudy (jdoe0001):

\(\large\bf 3^{\frac{1}{3}}=x\implies \sqrt[3]{3}=x\)

OpenStudy (anonymous):

where do I go from here now? \[Log_3(1/3)=x?\]

OpenStudy (anonymous):

I'm kind of lost on how we get to the end 3(1/3) part

OpenStudy (jdoe0001):

what part?

OpenStudy (anonymous):

After we get 1/3

OpenStudy (jdoe0001):

\(\large \bf {\color{red}{\cfrac{1}{3}}=log_\color{blue}{3}(x)\implies \color{blue}{3}^{\color{red}{\frac{1}{3}}}=x\implies \sqrt[3]{3}=x}\)

OpenStudy (jdoe0001):

all I did was convert it to exponent notation twice once for the outer function once for the inner one

OpenStudy (anonymous):

so its 3^1/3?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

and can be written at 3 root 3

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

gotcha so one more question if you don't mind me asking

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

Solve for X. 4 ^ (x- 4 ) = 3 Find an exact solution without decimals. You might want to use ln(x) or log(x).

OpenStudy (jdoe0001):

\(\bf 4^{x- 4} = 3\\ \quad \\ \textit{using log cancellation rule of }\quad \color{blue}{log_aa^x=x}\qquad thus\\ \quad \\ 4^{x- 4} = 3\implies log_4(4^{x- 4} )=log_4(3)\implies x-4=log_4(3)\\ \quad \\ \textit{using the log change of base rule of }\quad \color{blue}{log_ab=\cfrac{log_cb}{log_ca}}\qquad thus\\ \quad \\ x-4=log_4(3)\implies x-4=\cfrac{log_{10}3}{log_{10}4}\implies x=\cfrac{log_{10}3}{log_{10}4}+4\)

OpenStudy (anonymous):

how did we get the log_4(4^x-4)

OpenStudy (jdoe0001):

we take \(\bf log_4\) to both sides

OpenStudy (jdoe0001):

the idea being to use the log cancellation rule where the log base, MATCHES the base with the exponent

OpenStudy (anonymous):

hence why they are both log base 10

OpenStudy (jdoe0001):

why did I use the log change of base rule? well, technically I didn't have to but since calculators don't have \(\bf log_3 \ or \ log_4\ only \ log_{10} \ or \ ln\), thus

OpenStudy (anonymous):

so I typed in the answer and said I got it wrong :((

OpenStudy (jdoe0001):

now you know why the suggestion was "You might want to use ln(x) or log(x)"

OpenStudy (anonymous):

I typed log3/log4+4

OpenStudy (jdoe0001):

that's good, I used \(\bf log_{10}\) just to make it non-ambiguous, but yes, log with no base implies base 10

OpenStudy (jdoe0001):

ohhh hmmm, well... I assume the answer expected is a value, so you have to get the logs first and add 4 to that

OpenStudy (anonymous):

so just divide and add?

OpenStudy (jdoe0001):

sorry for my lag :( yes, you have to divide and add